Summarise the rules for writing of distribution of electrons in various shells for the first eighteen elements.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


The rules for writing of distribution of electrons in various shells for the first eighteen elements are:
(i) The maximum number of electrons present in a shell is formula—2n^2
n = orbit number i.e., 1, 2, 3
Maximum number of electrons in different shells are:
K shell n = 1 2n^2⇒ 2(1)^2= 2
L shell n = 2 2n^2⇒ 2(2)^2= 8
M shell n = 3 2n^2⇒ 2(3)^2= 18
N shell n = 4 2n^2⇒ 2(4)^2= 32
(ii) The maximum number of electrons that can be accommodated in the outermost orbit is 8.
(iii) Electrons are not accommodated in a given shell unless the inner shells are filled. (Shells are filled step-wise).


Thanks